your friend saved 12,000 sar during his studies in ksu. he wants to invest and he wants your help to choose one of the following two alternative investment opportunities. the first is buying 10 ounces of gold or buying 1,200 call options. the current price of gold is 1200 sar, and the 6- month european call option with a strike price 1185 sar is 12 sar. he strongly feels that the price of gold will increase. what is your advice to your friend (explain your advice)?
1
Expert's answer
2020-02-24T09:46:00-0500


If the price of gold rises to 1300 in 6 months:


total option price=1185+12=1197

gold price 1300

1300-1197=103


Income from gold is greater than from an option after 6 months, therefore it is more profitable to invest in gold